# Spontaneous orbital selective Mott phase in the two band Hubbard model

### v1

@@ Abstract
Quantum materials featuring both itinerant and localized degrees of freedom exhibit numerous exotic phases and transitions that deviate from the Ginzburg-Landau paradigm. This work uses the composite operator formalism to examine two-orbital strongly correlated Hubbard models. We observe the spontaneous breaking of orbital symmetry, where the electron density in one of the orbitals reaches half-filling, resulting in an orbitally selective Mott phase (OSMP). This broken symmetry phase becomes unstable at a critical average electronic density away from half-filling. Furthermore, significant orbital differentiation persists up to a moderate inter-orbital hopping, beyond which the system abruptly transitions to an orbitally uniform phase. In the OSMP phase, the electrons in the two orbitals are weakly hybridized, resulting in a small Fermi surface. The volume of the Fermi surface jumps at the transition from the OSMP to the orbitally uniform phase. We also discuss the physical mechanisms leading to the collapse of the OSMP phase under different perturbations.
@@ I Introduction
Orbital degrees of freedom can lead to intriguing phenomena in strongly correlated quantum matter such as spontaneous orbital symmetry breaking, Hund metallicity [12]. The orbital selective Mott phase (OSMP) was first proposed to explain the coexistence of metallic and magnetic properties in ruthenates [13, 14] and has since been recognized as key for understanding the normal [15] and superconducting properties [16, 17] of iron-based superconductors. It is also crucial for understanding the local versus nonlocal duality in $f$-electron-based compounds [18]. The OSMP is characterized by the selective localization of electrons in specific orbitals, while electrons in other orbitals remain itinerant. Initial evidence came from Angle-resolved photoemission spectroscopy (ARPES) studies [19] in $\mathrm{Ca_{1.8}Sr_{0.2}RuO_{4}}$, followed by similar observations in iron-chalcogenides [20]. In these materials, the spectral weight of the $d_{xy}$ orbitals vanishes upon cooling, while the spectral weight of other $3d$ orbitals remains non-zero. More recently, theoretical proposals suggest that OSMP could also emerge in twisted bilayer systems such as twisted dichalcogenides [21] and twisted trilayer graphene [22].
@@
The OSMP phase can be defined as the complete localization of one orbital entirely decoupled from the other metallic orbitals [26]. The stability of the OSMP phase with interorbital hoppings remains a subject of debate. While there is agreement among various methods that an OSMP exists when the hopping between different orbitals is negligible, the situation becomes more intricate when it is non-zero. The slave-spin formalism indicates the stability of the OSMP phase with inter-orbital hopping [27, 28], with the hybridization of the localized orbitals to the others renormalized to zero. Single-site dynamical mean field theory (DMFT) claims [26] that any inter-orbital hopping should lead to a finite hybridization. Nevertheless, both methods agree that even with finite hybridization, orbital differentiation exists with a significant difference in effective mass.
@@
This paper investigates the degenerate two-band Hubbard model, also known as the bilayer Hubbard model [29], with only intra-orbital on-site repulsion $U$. We examine the effect of finite intraorbital hopping while varying the total electron density in the strong correlation regime. In this regime, it is suitable to start from the atomic Hamiltonian $(t=0)$ and generate a set of quasiparticle operators for the ground state. In these schemes, electrons in a strongly interacting regime are generally considered a sum of weakly interacting excitations [30]. These excitations correspond to holons and doublons in the strongly correlated limit of the one-band Hubbard model, both of which have identical quantum numbers to the electron. This results in the emergence of the lower and upper Hubbard bands, characterized predominantly by holons and doublons, respectively. Such schemes can be contrasted with slave-boson methods [31, 32, 33] where the electron fractionalizes into emergent operators that carry fractions of the electron’s charge and spin. This introduces gauge freedom, making the model equivalent to solving gauge theories.
@@
Using the composite operator method (COM) [34, 35, 36, 37, 38], we explore the strongly correlated limit of the two band Hubbard model. This technique has previously been applied to specific solutions of the two-band Hubbard model [39], demonstrating the existence of an OSMP phase. The objective of this paper is to perform a detailed study of the physical properties and orbital symmetry-breaking phase in such a model. Our central finding is the emergence of a spontaneous orbital differentiation phase for moderate values of interorbital hopping without explicit symmetry-breaking parameters between the two orbitals [29]. Our results offer a complementary perspective to slave-boson and DMFT approaches in similar models and provide novel insights into the impact of finite inter-orbital hopping on the OSMP phase and its abrupt destabilization. Additionally, we investigate the correspondence between the Kondo lattice model and this two-orbital Hubbard model [40, 41], particularly examining whether complete Kondo breakdown [42] occurs in the OSMP phase.
@@
The plan for the rest of the paper is as follows. The next section discusses the model and the composite operator formalism for the two-band Hubbard model. The detailed expressions for self-consistency equations and physical observables are presented in Appendix LABEL:Appendix:Details_COM. We present the overall phase diagram for the strength of fixed interactions in Sec. (III.1). The subsequent subsections discuss the bands, the Fermi surface, the density of states, and quasiparticle residues identifying the two phases. We conclude with a discussion of the results in Sec. (IV).
@@ II Model and method
Here, $c_{i\alpha\sigma}$ $(c^{\dagger}_{i\alpha\sigma})$ annihilates (creates) an electron at site $i$ with orbital $\alpha$ and spin $\sigma$, where $\sigma=\uparrow,\downarrow$ for spin-$1/2$ electrons and ${\alpha=x,y}$ in a square lattice. The number operator is defined as $\hat{n}_{i\alpha\sigma}=c^{\dagger}_{i\alpha\sigma}c_{i\alpha\sigma}$, and the local electron density is the expectation value of the number operator, $n_{i\alpha\sigma}=\langle\hat{n}_{i\alpha\sigma}\rangle$.
The first term represents intraorbital hopping between neighboring sites $i$ and $j$, set to $t_{ij}=t$ if $i$ and $j$ are nearest neighbors; otherwise, zero. We set $t^{x}=t^{y}=t$ in this work, and all our energy scales are in units of $t=1$. The second term describes on-site interorbital hopping with a coupling constant $\lambda$. The third term accounts for intra-orbital on-site repulsion between electrons. We assume that this intra-orbital repulsion has an equal strength for both orbitals $U^{x}=U^{y}=U>0$. The chemical potential $\mu$ fixes the total electron density of the system $n=(1/N)\sum_{i,\alpha,\sigma}\langle\hat{n}_{i\alpha\sigma}\rangle$ which can vary from $n\in[0,4]$. Here is the number of lattice points in the system. Our analysis of the system is valid in the strong correlation regime where $U\gg t,\lambda$.
@@ II.1 Composite operators - Equation of Motion
In the strong correlation regime, the excitation of the singly occupied ground state are given by holon ($\xi$) and doublon ($\eta$) single orbital operators
@@
The composite operator method (COM) [37] is an equation of motion technique to compute the electronic Green’s function $\mathcal{G}(\mathbf{k},\omega)$ and hence other physical quantities. For the two-band Hubbard model with only intra-orbital electronic repulsion, the paramagnetic composite operator basis is defined by
@@
where we defined the current as ${\mathbf{j}(\tau)=-\left[\mathcal{H},\mathbf{\Psi}\right](\tau)}$. Next we define the normalization matrix $\mathds{I}$ and $\mathds{M}$-matrix as
$\mathds{I}$ =$=\left\langle\{\Psi(0),\Psi^{\dagger}(0)\}\right\rangle,$ (7)
$\mathds{M}(\tau)$ $=\left\langle T_{\tau}\left(\mathbf{j}(\tau)\mathbf{\Psi}(0)\right)\right\rangle.$ (8)
@@
$\mathbf{j}(\tau)\approx\mathds{E}\Psi(\tau).$ (9)
@@
$\mathds{M}(0)$ =$=\left\langle\{\mathbf{j}(0),\Psi^{\dagger}(0)\}\right\rangle,$ (10)
@@ II.2 Computation of E and Matrix
Here $\gamma(\mathbf{k})=2t\left(\cos(k_{x})+\cos(k_{y})\right)$. Moreover, in $\mathds{M}^{\prime}$, the orbital indices $x$ and $y$ are interchanged compared to $\mathds{M}$. For instance, $\mathds{M}_{33}$ has the same expression as $\mathds{M}_{11}$ with all the subscript $x$ and $y$ interchanged. We also define average orbital electron density as
@@ II.3 Correlation function and self-consistency
Using the on-site ($C^{ab}_{0}$) and nearest neighbor ($C^{ab}_{1}$) correlation functions we can write the orbital density as
@@
We define the difference in orbital density as ${\delta n=(n_{x}-n_{y})/2}$ as a self-consistent parameter, and fix the total density $n=n_{x}+n_{y}$ using the chemical potential $\mu$. Similarly the $e$-parameters can be recalculated by
@@
However, $p_{\alpha\beta}$ is a two-particle correlation function that cannot be evaluated directly from single-particle correlation functions. For the one-band Hubbard model, $p$ has also been computed by enforcing the Pauli principle $\langle\xi_{i}\eta_{i}^{\dagger}\rangle=0$ [43]. In the two-band Hubbard model, such approach leaves the computation of interorbital $p_{xy}$ ambiguous. Therefore, we evaluate $p_{\alpha\beta}$ for $\alpha,\beta\in\{x,y\}$ using the Roth decoupling scheme [44], which we outline in Appendix (A). The $p$-parameter depends on three distinct nearest neighbor correlation functions as seen in Eq. (23). The expression for density-density correlations in terms of $C^{ab}_{0}$ and $C^{ab}_{1}$ is given in Eq. (45). Similarly the expression for $\langle S^{-}_{i\alpha}S^{+}_{j\beta}\rangle$ in terms of correlation functions is presented in Eq. (46) and $\langle\Delta_{i\alpha}\Delta^{\dagger}_{j\beta}\rangle$ is presented in Eq. (47). Using these one can recalculate two intraorbital and one interorbital $p_{\alpha\beta}$.
Subsequently, we solve these nine coupled self-consistency equations. In the orbital-selective Mott phase, one of the orbitals becomes half-filled. Therefore, we initialize with the guess $\delta n=(2-n)$ and solve the self-consistency equations. If the converged solution has one of the orbital densities approaching half-filling $n_{\alpha}\rightarrow 1$, then that solution represents the OSMP phase. In the literature [45, 46, 26], the OSMP phase is defined by the vanishing hybridization between the localized electrons in a given orbital and the other itinerant orbitals. If the localized orbital has a density very close to, but not exactly, $1$ due to finite hybridization, then the system’s physical properties are effectively identical to those of a conventional OSMP phase for all practical purposes. In the following, we will extend the term OSMP to include such phases. Conversely, a solution with $\delta n=0$ represents an orbitally uniform phase (OUP). We compare the energy per site $E_{s}$ of these two solutions to identify the lower energy solution.
@@ III Results
We perform the calculations for a square lattice with linear dimension $L=300$ with periodic boundary conditions. The repulsion strength is fixed at $U=20t$ for the results presented in the main text. However, we have ensured the reliability of our qualitative conclusions across various repulsion strengths ranging from $U=9t$ to $U=20t$. Our investigation spans a range of average electronic densities between $n=1.50$ and $n=1.94$. We focus on discerning orbital symmetry breaking within systems invariant under spin-rotation and translation.
@@ III.1 Phase diagram
Figure 1: (A) Displays the phase diagram as a function of electron density $n$ and inter-orbital hopping $\lambda$ obtained for $U=20t$. Across all parameters, an orbitally uniform phase state exists, characterized by equal electron densities in both orbitals. However, an orbital selective Mott phase emerges above the red line, with one orbital spontaneously reaching half-filling. (B) Depicts the disparity in orbital densities ($\delta n$) as a function of $\lambda$ for multiple average electron densities ($n$), with colors matching those in (A). For small $\lambda/t$, one orbital closely approaches half-filling before dropping to equal density, indicating a sharp transition. (C) Compares the energy per site between the two solutions. The thick lines are for the OSMP solution, with the dashed line corresponding to the OUP one. The green area in (A) denotes where the OSMP has lower energy than the OUP solution. The region between black and red lines in (A) has a metastable OSMP phase.
@@
In Fig. (1A), we present the phase diagram of the model in the total electron density $n$ and inter-orbital hopping $\lambda$ plane, with $U=20t$ for each orbital. The orbitally uniform phase exhibits identical orbital electron densities, akin to the scenario described in ref.[38] for disconnected orbitals. Such a uniform orbital density solution is found across the entire phase diagram. However, an orbital selective Mott phase emerges above the black line. In the OSMP phase, the density of one orbital spontaneously approaches half-filling.
In Fig. (1B) illustrates the difference in electron density between the two orbitals $|\delta n|$. Each line color corresponds to a specific total electron density associated with colored dashed lines in Fig. (1A). Consequently, as $\lambda$ exceeds a critical value, $\delta n$ diminishes to zero, indicating a sharp transition to a state with equal density at each orbital. At $\lambda=0$, the electron density at one of the orbital is strictly at half-filling. We can gain insight into this solution by comparing it with the one-band Hubbard model. When one orbital is half-filled, the chemical potential $\mu$ can be adjusted to ensure the self-consistent equations are met for the other orbital. However, if $\mu$ evolved such that the half-filled band transitions to a metallic state, the self-consistent conditions for the previously half-filled orbitals are no longer satisfied, leading to the disappearance of the OSMP phase. When $\lambda\neq 0$, before transitioning to the orbitally uniform phase phase, one orbital maintains a density close to half-filling for finite interorbital hoppings. The critical $\lambda$ obtained from this analysis shows that the self-consistent OSMP solutions vanishes as the interorbital hopping increases.
@@
However, the stability of the OSMP phase is determined by comparing the energy per site of the OSMP and the OUP phases. In Fig.1C), we show the energy per site for both states as a function of $\lambda$ across various electron densities, using the same color code as in Fig.(1B). The dashed line represents the OUP phase, while the continuous line depicts the OSMP. The OSMP phase becomes the ground state when its energy is lower than that of the orbitally uniform phase. The OSMP phase is energetically favorable near half-filling ($n=2$) and for weak to moderate interorbital hopping. The colored area of Fig.(1A) indicates the parameter range where the OSMP is the lower energy phase. Although the OSMP can stabilize below the thick black lines into the white region, it exhibits higher energy than the orbitally uniform phase. Only OUP solutions are observed below the thick red lines in Fig. (1A).
@@
We can understand the stability of OSMP phase in the absence of interorbital hopping by studying the energy per site $E_{s}$ of the one-band Hubbard model as a function of $n$. See Appendix (C) for details. For $\lambda=0$, OMSP is favored phase whenever $E_{s}(1)+E_{s}(1-2\delta)<2E_{s}(1-\delta)$ with $\delta$ the doping. We checked that this is satisfied only close to half-filling indicating the regime of stability of the OSMP phase.
@@
Consequently, we observe a spontaneous symmetry breaking of the orbital degree of freedom in the two-band Hubbard model. In the OSMP, one of the orbitals approaches the half-filled Mott state. As expected, such a phase is stabilized when the total electron density is close to half-filling  [^1] Since at $n=2$, both the orbitals become Mott insulating.. Additionally, the OSMP phase persists to a critical interorbital hopping $\lambda_{c}$, indicating a sharp transition. Such critical $\lambda_{c}$ increases as we approach the particle-hole symmetric point. In Sec. (III.6), we explore the dependence of the system for varying interaction strength $U$. Subsequent subsections characterize the two phases using multiple physical observables.
III.2 Bands and interorbital hybridization

### v2

# Spontaneous layer selective Mott phase in the bilayer Hubbard model

@@ Abstract
Quantum materials featuring both itinerant and localized degrees of freedom exhibit numerous exotic phases and transitions that deviate from the Ginzburg-Landau paradigm. This work uses the composite operator formalism to examine the bilayer strongly correlated Hubbard model. We observe the spontaneous breaking of layer symmetry, where the electron density in one of the layer reaches half-filling, resulting in a layer selective Mott phase (LSMP). This broken symmetry phase becomes unstable at a critical average electronic density away from half-filling. Furthermore, significant layer differentiation persists up to a moderate inter-layer hopping, beyond which the system abruptly transitions to an layer uniform phase (LUP). In the LSMP phase, the electrons in the two layers are weakly hybridized, resulting in a small Fermi surface. The volume of the Fermi surface jumps at the transition from the LSMP to the uniform phase. We also discuss the physical mechanisms leading to the collapse of the LSMP phase under different perturbations.
@@ I Introduction
Orbital degrees of freedom can lead to intriguing phenomena in strongly correlated quantum matter such as spontaneous orbital symmetry breaking, Hund metallicity [12]. The orbital selective Mott phase (OMSP) was first proposed to explain the coexistence of metallic and magnetic properties in ruthenates [13, 14] and has since been recognized as key for understanding the normal [15] and superconducting properties [16, 17] of iron-based superconductors. It is also crucial for understanding the local versus nonlocal duality in $f$-electron-based compounds [18]. The OSMP is characterized by the selective localization of electrons in specific orbitals, while electrons in other orbitals remain itinerant. Initial evidence came from Angle-resolved photoemission spectroscopy (ARPES) studies [19] in $\mathrm{Ca_{1.8}Sr_{0.2}RuO_{4}}$, followed by similar observations in iron-chalcogenides [20]. In these materials, the spectral weight of the $d_{xy}$ orbitals vanishes upon cooling, while the spectral weight of other $3d$ orbitals remains non-zero. More recently, theoretical proposals suggest that OSMPlike phase could also emerge in twisted bilayer systems such as twisted dichalcogenides [21] and twisted trilayer graphene [22].
@@
The orbital selective phase can be defined as the complete localization of one orbital entirely decoupled from the other metallic orbitals [26]. The stability of the OSMP phase with interorbital hoppings remains a subject of debate. While there is agreement among various methods that an OSMP exists when the hopping between different orbitals is negligible, the situation becomes more intricate when it is non-zero. The slave-spin formalism indicates the stability of the OSMP phase with inter-orbital hopping [27, 28], with the hybridization of the localized orbitals to the others renormalized to zero. Single-site dynamical mean field theory (DMFT) claims [26] that any inter-orbital hopping should lead to a finite hybridization. Nevertheless, both methods agree that even with finite hybridization, orbital differentiation exists with a significant difference in effective mass.
@@
This paper investigates the degenerate bilayer Hubbard model [29, 30, 31, 32, 33], with only on-site repulsion $U$. Since we do not consider Hund’s coupling necessary for stabilizing the OSMP phase [24, 25], we refer to the resulting symmetry broken phase as layer selective Mott phase (LSMP) to distinguish it from the OSMP. Such layer-selective Mott phases can potentially be important for delafossites [34] and nickelates [35] and other layered systems. We examine the effect of finite inter-layer hopping while varying the total electron density in the strong correlation regime. In this regime, it is suitable to start from the atomic Hamiltonian $(t=0)$ and generate a set of quasiparticle operators for the ground state. In these schemes, electrons in a strongly interacting regime are generally considered a sum of weakly interacting excitations [36]. These excitations correspond to holons and doublons in the strongly correlated limit of the one-band Hubbard model, both of which have identical quantum numbers to the electron. This results in the emergence of the lower and upper Hubbard bands, characterized predominantly by holons and doublons, respectively. Such schemes can be contrasted with slave-boson methods [37, 38, 39] where the electron fractionalizes into emergent operators that carry fractions of the electron’s charge and spin. This introduces gauge freedom, making the model equivalent to solving gauge theories.
@@
Using the composite operator method (COM) [40, 41, 42, 43, 44, 45], we explore the strongly correlated limit of the two layer Hubbard model. This technique has previously been applied to specific solutions of the two-band Hubbard model [46], demonstrating the existence of a selective Mott phase. The objective of this paper is to perform a detailed study of the physical properties and layer symmetry-breaking phase in such a model. Our central finding is the emergence of a spontaneous layer differentiation phase for moderate values of interlayer hopping without explicit symmetry-breaking parameters between the two layers [29]. Our results offer a complementary perspective to slave-boson and DMFT approaches in similar models and provide novel insights into the impact of finite inter-layer hopping on the LSMP phase and its abrupt destabilization. Additionally, we investigate the correspondence between the Kondo lattice model and bilayer Hubbard model [47, 48], particularly examining whether complete Kondo breakdown [49] occurs in the LSMP phase.
@@
The plan for the rest of the paper is as follows. The next section discusses the model and the composite operator formalism for the two-band Hubbard model. The detailed expressions for self-consistency equations and physical observables are presented in Appendix (B). We present the overall phase diagram for the strength of fixed interactions in Sec. (III.1). The subsequent subsections discuss the bands, the Fermi surface, the density of states, and quasiparticle residues identifying the two phases. We conclude with a discussion of the results in Sec. (IV).
@@ II Model and method
Here, $c_{i\alpha\sigma}$ $(c^{\dagger}_{i\alpha\sigma})$ annihilates (creates) an electron at site $i$ with layer $\alpha$ and spin $\sigma$, where $\sigma=\uparrow,\downarrow$ for spin-$1/2$ electrons and ${\alpha=x,y}$ in a square lattice. The number operator is defined as $\hat{n}_{i\alpha\sigma}=c^{\dagger}_{i\alpha\sigma}c_{i\alpha\sigma}$, and the local electron density is the expectation value of the number operator, $n_{i\alpha\sigma}=\langle\hat{n}_{i\alpha\sigma}\rangle$.
The first term represents intra-layer hopping between neighboring sites $i$ and $j$, set to $t_{ij}=t$ if $i$ and $j$ are nearest neighbors; otherwise, zero. We set $t^{x}=t^{y}=t$ in this work, and all our energy scales are in units of $t=1$. The second term describes on-site inter-layer hopping with a coupling constant $\lambda$. The third term accounts for intra-layer on-site repulsion between electrons. We assume that this intra-layer repulsion has an equal strength for both layers $U^{x}=U^{y}=U>0$. The chemical potential $\mu$ fixes the total electron density of the system $n=(1/N)\sum_{i,\alpha,\sigma}\langle\hat{n}_{i\alpha\sigma}\rangle$ which can vary from $n\in[0,4]$. Here is the number of lattice points in the system. Our analysis of the system is valid in the strong correlation regime where $U\gg t,\lambda$.
@@ II.1 Composite operators - Equation of Motion
In the strong correlation regime, the excitation of the singly occupied ground state are given by holon ($\xi$) and doublon ($\eta$) single layer operators
@@
Since we are focusing on an interacting model with only on-site, intra-layer interactions, the $\xi$ and $\eta$ operators represent the most relevant excitations near half-filling. To capture the physics of more complex models, such as those including significant inter-layer Coulomb interactions, the operator basis must be expanded. This is the case in multi-orbital models with a Hund term, where additional bands can be observed [50, 51]. In contrast, our approximation will always yield four bands for two layers. The composite operator method (COM) [43] is an equation of motion technique to compute the electronic two-point Green’s function $\mathcal{G}(\mathbf{k},\omega)$ and hence other physical quantities. For the two-band Hubbard model with only intra-layer electronic repulsion, the paramagnetic composite operator basis is defined by
@@
where we defined the current as ${\mathbf{j}(\tau)=\left[\mathcal{H},\mathbf{\Psi}\right](\tau)}$. Next we define the normalization matrix $\mathds{I}$ and $\mathds{M}$-matrix as
$\mathds{I}$ =$=\left\langle\{\mathbf{\Psi}(0),\mathbf{\Psi}^{\dagger}(0)\}\right\rangle,$ (7)
$\mathds{M}(\tau)$ $=\left\langle T_{\tau}\left(\mathbf{j}(\tau)\mathbf{\Psi}^{\dagger}(0)\right)\right\rangle.$ (8)
@@
$\mathbf{j}(\tau)\approx\mathds{E}\mathbf{\Psi}(\tau).$ (9)
@@
$\mathds{M}(0)$ =$=\left\langle\{\mathbf{j}(0),\mathbf{\Psi}^{\dagger}(0)\}\right\rangle,$ (10)
@@ II.2 Computation of E and Matrix
Here $\gamma(\mathbf{k})=2t\left(\cos(k_{x})+\cos(k_{y})\right)$. Moreover, in $\mathds{M}^{\prime}$, the layer indices $x$ and $y$ are interchanged compared to $\mathds{M}$. For instance, $\mathds{M}_{33}$ has the same expression as $\mathds{M}_{11}$ with all the subscript $x$ and $y$ interchanged. We also define average layer electron density as
@@ II.3 Correlation function and self-consistency
Using the on-site ($C^{ab}_{0}$) and nearest neighbor ($C^{ab}_{1}$) correlation functions we can write the layer density as
@@
We define the difference in layer density as ${\delta n=(n_{x}-n_{y})/2}$ as a self-consistent parameter, and fix the total density $n=n_{x}+n_{y}$ using the chemical potential $\mu$. Similarly the $e$-parameters can be recalculated by
@@
However, $p_{\alpha\beta}$ is a two-particle correlation function that cannot be evaluated directly from single-particle correlation functions. For the one-band Hubbard model, $p$ has also been computed by enforcing the Pauli principle $\langle\xi_{i}\eta_{i}^{\dagger}\rangle=0$ [52, 53]. In the two-band Hubbard model, such approach leaves the computation of interlayer $p_{xy}$ ambiguous. Therefore, we evaluate $p_{\alpha\beta}$ for $\alpha,\beta\in\{x,y\}$ using the Roth decoupling scheme [54], which we outline in Appendix (A). The $p$-parameter depends on three distinct nearest neighbor correlation functions as seen in Eq. (23). The expression for density-density correlations in terms of $C^{ab}_{0}$ and $C^{ab}_{1}$ is given in Eq. (46). Similarly the expression for $\langle S^{-}_{i\alpha}S^{+}_{j\beta}\rangle$ in terms of correlation functions is presented in Eq. (47) and $\langle\Delta_{i\alpha}\Delta^{\dagger}_{j\beta}\rangle$ is presented in Eq. (48). Using these one can recalculate two intra-layer and one inter-layer $p_{\alpha\beta}$.
Subsequently, we solve these nine coupled self-consistency equations. In the layer-selective Mott phase, one of the layers becomes half-filled. Therefore, we initialize with the guess $\delta n=(2-n)$ and solve the self-consistency equations. If the converged solution has one of the layer densities approaching half-filling $n_{\alpha}\rightarrow 1$, then that solution represents the LSMP phase. In the literature [55, 56, 26], the LSMP phase is defined by the vanishing hybridization between the localized electrons in a given layer and the other itinerant layers. If the localized layer has a density very close to, but not exactly, $1$ due to finite hybridization, then the system’s physical properties are effectively identical to those of a conventional LSMP phase for all practical purposes. In the following, we will extend the term LSMP to include such phases. Conversely, a solution with $\delta n=0$ represents a uniform phase (LUP). We compare the energy per site $E_{s}$ of these two solutions to identify the lower energy solution.
@@ III Results
We perform the calculations for a square lattice with linear dimension $L=300$ with periodic boundary conditions. The repulsion strength is fixed at $U=20t$ for the results presented in the main text. However, we have ensured the reliability of our qualitative conclusions across various repulsion strengths ranging from $U=9t$ to $U=20t$. Our investigation spans a range of average electronic densities between $n=1.50$ and $n=1.94$. We focus on discerning layer symmetry breaking within systems invariant under spin-rotation and translation.
@@ III.1 Phase diagram
Figure 1: (A) Displays the phase diagram as a function of electron density $n$ and inter-layer hopping $\lambda$ obtained for $U=20t$. Across all parameters, a layer uniform phase state exists, characterized by equal electron densities in both layers. However, a layer selective Mott phase emerges above the red line, with one layer spontaneously reaching half-filling. (B) Depicts the disparity in layer densities ($\delta n$) as a function of $\lambda$ for multiple average electron densities ($n$), with colors matching those in (A). For small $\lambda/t$, one layer closely approaches half-filling before dropping to equal density, indicating a sharp transition. (C) Compares the energy per site between the two solutions. The dashed lines are for the LUP solution, with the thick line corresponding to the LSMP one. The green area in (A) denotes where the LSMP has lower energy than the LUP solution. The region between black and red lines in (A) has a metastable LSMP phase.
@@
In Fig. (1A), we present the phase diagram of the model in the total electron density $n$ and inter-layer hopping $\lambda$ plane, with $U=20t$ for each layer. The uniform phase exhibits identical layer electron densities, akin to the scenario described in ref.[45] for disconnected layers. Such a uniform layer density solution is found across the entire phase diagram. However, a layer selective Mott phase emerges above the black line. In the LSMP phase, the density of one layer spontaneously approaches half-filling.
In Fig. (1B) illustrates the difference in electron density between the two layers $|\delta n|$. Each line color corresponds to a specific total electron density associated with colored dashed lines in Fig. (1A). Consequently, as $\lambda$ exceeds a critical value, $\delta n$ diminishes to zero, indicating a sharp transition to a state with equal density at each layer.
@@
However, the stability of the LSMP phase is determined by comparing the energy per site of the LSMP and the LUP phases. In Fig.1C), we show the energy per site for both states as a function of $\lambda$ across various electron densities, using the same color code as in Fig.(1B). The dashed line represents the LUP phase, while the continuous line depicts the LSMP. The LSMP phase becomes the ground state when its energy is lower than that of the uniform phase. The LSMP phase is energetically favorable near half-filling ($n=2$) and for weak to moderate inter-layer hopping. The colored area of Fig.(1A) indicates the parameter range where the LSMP is the lower energy phase. Although the LSMP can stabilize below the thick black lines into the white region, it exhibits higher energy than the uniform phase. Only LUP solutions are observed below the thick red lines in Fig. (1A).
We can understand the origin of LSMP phase in the absence of inter-layer hopping i.e. $\lambda=0$, where the two layers are disconnected. Note that the electron density in the two layers is constrained by the total electron density, i.e., $n=n_{x}+n_{y}$, which is determined by the same chemical potential, $\mu$. In the uniform phase, both layers share the same chemical potential which leads to the uniform solution of the single-layer Hubbard system with an electronic density of $n/2$ [45].
An alternative scenario arises when one layer becomes fully occupied spontaneously and falls entirely below the chemical potential for a given total electron density. To satisfy the density constraint, the chemical potential of the second layer adjusts to achieve a filling of $n_{y}=n-1$. As the total electron density decreases, the chemical potential is reduced accordingly to adjust the filling of the metallic layer, although the band structure of the filled band remains unchanged. Once the chemical potential touches the top of the half-filled band, the system transitions to an uniform phase.
Figure 2: Energy per site for the single-band Hubbard model as a function of electron density $n$. The red diamond, green square, and orange circle indicate the density and energy at $n=1.0$, $n=0.9$, and $n=0.8$ respectively. For two layers, if the the total electron density is $n=1.8$. Without interlayer hopping this demonstrates that it is energetically favorable to have one layer at $n=1$ and the other at $n=0.8$ (LSMP phase), rather than both layers at $n=0.9$ (OUP phase).
Previous studies using the paramagnetic composite operator method for the single-band Hubbard model have observed negative compressibility near half-filling [42, 40] and a small energy dip near half-filling in other models [57], indicating that the uniform solution is unstable towards phase separation. In this work, phase separation occurs in the layer degrees of freedom, with one layer becoming half-filled while the other accommodates all the doped holes. Consequently, phase separation is the key mechanism driving the selective Mott phase in the paramagnetic limit. It remains unclear whether this observation is an artifact of the Hubbard operator method. Advancing the method to the next order, as suggested in Ref. [58], can provide insight into these questions.
@@
The stability of the LSMP phase at $\lambda=0$ can be understood by studying the energy per site $E_{s}$ of the one-band Hubbard model as a function of $n$. For $\lambda=0$, LSMP is favored phase whenever $E_{s}(1)+E_{s}(1-2\delta)<2E_{s}(1-\delta)$ with $\delta$ the doping. We checked that this is satisfied only close to half-filling indicating the regime of stability of the LSMP phase.
@@
Consequently, we observe a spontaneous symmetry breaking of the layer degree of freedom in the bilayer Hubbard model. In the LSMP, one of the layers approaches the half-filled Mott state. As expected, such a phase is stabilized when the total electron density is close to half-filling  [^1] Since at $n=2$, both the layers become Mott insulating.. Additionally, the LSMP phase persists to a critical inter-layer hopping $\lambda_{c}$, indicating a sharp transition. Such critical $\lambda_{c}$ increases as we approach the particle-hole symmetric point. In Sec. (III.6), we explore the dependence of the system for varying interaction strength $U$. Subsequent subsections characterize the two phases using multiple physical observables.
In this section, we presented results for on-site inter-layer hopping. In App. C, we demonstrate that the results remain qualitatively similar with different form of inter-layer hopping like extended $s$-wave and $d$-wave hopping.
III.2 Bands and inter-layer hybridization